# Designing metasurface optical interfaces for solid-state qubits using many-body adjoint shape optimization

### v1

@@ IV Discussion and Outlook
The goal of the optimized structure is to maximize photon collection from a 1-µm-deep NV center directly into an optical fiber, obviating the need for high-NA free-space optics. We are not aware of any superior solution for this task. We simulate an overall transmission efficiency of 3.61% (including all emitted broadband photons assuming a (100) diamond surface) into the collection aperture of a multi-mode fiber. This is a factor of 3.5 improvement over a fiber-coupled SIL or a factor of 26 improvement over a fiber-coupled flat air-diamond interface; a SIL can only reach its peak efficiency when combined with a high-NA optic and remains the best-performing option in that case. To predict the device performance in an experimental setting, we compare to Ref. Huang et al. (2019), in which the authors measured a saturation photon count rate of 87.3 $\pm$ 2.8 kcps (kilocounts per second) from an NV center imaged through a similar metasurface coupled to a multi-mode fiber. In simulations, the optimized metasurface outperforms that structure by roughly 60%, and we would expect to obtain a photon count rate around 150 kcps under otherwise identical experimental conditions and with similar losses from fabrication errors. Additionally, the metasurface achieves this greater performance while being much more compact (5 µm vs 28 µm diameter) and having a shorter depth of focus (1 µm vs 20 µm), which allows for easier alignment and fabrication.
@@
The many-body shape optimization methodology we present achieves comparable performance to the more common adjoint-based technique of topology optimization. While topology optimization enforces its fabrication constraints by subsequently applying filters over multiple optimization phases, MBSO allows for constraint functions to be defined explicitly from the parameterization and applied throughout a single-phase optimization. As a result, the shape-optimized structure converges much faster, using less than half as many sets of simulations as a topology-optimized structure (200 vs 471) — the computational requirements of both types of adjoint optimization are primarily limited by the number of simulations required. The shape-optimized structure also required an order of magnitude fewer parameters than the topology-optimized structure (1280 vs 40,401); although this did not strongly impact the computational performance in this case, it could result in significantly faster gradient calculations for larger-scale metasurfaces.
@@
As an adjoint-based optimization strategy, MBSO’s scalability to larger geometrical areas is primarily limited by the computational requirements in memory and the time to perform the full-wave simulations at each iteration. Several techniques may be used to increase the viability of large-scale optimizations. Computational advancements such as GPU acceleration can enable larger-scale full-wave simulations Warren et al. (2019). Past authors have performed shape optimizations on structures an order of magnitude larger than the present work by applying symmetrical boundary conditions and using rectangular elements aligned to the FDTD mesh Mansouree et al. (2021) or by using Maxwell solvers based on boundary integral methods Sideris et al. (2019); Garza and Sideris (2022) — both strategies are particularly well-suited to shape optimizations. Other works have avoided full-wave simulations by using local periodic approximations Pestourie et al. (2018) and coupled mode theory Zhou et al. (2021) in limited regimes or deep learning methods to predict forward and inverse problems Zhelyeznyakov et al. (2021); Kanmaz et al. (2023); Jiang and Fan (2019). The MBSO parameterization method could easily be adapted into any of these alternative optimization strategies.
@@
The MBSO methodology is readily applicable to the inverse design a variety of structures, especially when there is a need for specific shapes or distinct elements. Broadly, it can be used in lieu of topology optimization for any inverse design problem. In particular, MBSO is naturally suited for the inverse design of photonic structures such as photonic crystals, large-area metasurfaces, and diffraction gratings that are typically composed of arrays of simple shapes. Additionally, MBSO could be useful for designing reconfigurable metasurfaces — which have potential applications including imaging, beam steering for LiDAR, and displays —  such that the design space is limited to a fixed number of meta-atoms that can be independently addressed by some tuning mechanism Gu et al. (2022). For certain materials, it may be challenging to fabricate arbitrary shapes due to facet-selective etching or growth. In these instances, MBSO could be easily adapted to keep any edges aligned to particular crystal facets in order to optimize structures exactly how they would be fabricated, potentially enabling the inverse design of photonic structures onto new material platforms.
@@ Acknowledgements.
This work was supported by the National Science Foundation under awards ECCS-1842655 (A.R.K., L.C.B.) and ECCS-2129183 (A.R.K., L.C.B.), by the National Science Foundation Graduate Research Fellowship Program, Grant No. DGE-1845298 (A.R.K.), and by the Air Force Office of Scientific Research Multidisciplinary University Research Initiative under grant FA9550-21-1-0312 (N.E.). The authors thank Brian Edwards for fruitful conversations and Mathieu Ouellet for assistance with the three-dimensional image in Fig. 1(a).

### v2

@@ IV Discussion and Outlook
The goal of the optimized structure is to maximize photon collection from a 1-µm-deep NV center directly into an optical fiber, obviating the need for high-NA free-space optics. We are not aware of any superior solution for this task. We simulate an overall transmission efficiency of 3.61% (including all emitted broadband photons assuming a (100) diamond surface) into the collection aperture of a multi-mode fiber. This is a factor of 3.5 improvement over a fiber-coupled SIL or a factor of 26 improvement over a fiber-coupled flat air-diamond interface; a SIL can only reach its peak efficiency when combined with a high-NA optic and remains the best-performing option in that case. To predict the device performance in an experimental setting, we compare to Ref. [20], in which the authors measured a saturation photon count rate of 87.3 $\pm$ 2.8 kcps (kilocounts per second) from an NV center imaged through a similar metasurface coupled to a multi-mode fiber. In simulations, the optimized metasurface outperforms that structure by roughly 60%, and we would expect to obtain a photon count rate around 150 kcps under otherwise identical experimental conditions and with similar losses from fabrication errors. Additionally, the metasurface achieves this greater performance while being much more compact (5 µm vs 28 µm diameter) and having a shorter depth of focus (1 µm vs 20 µm), which allows for easier alignment and fabrication.
@@
The many-body shape optimization methodology we present achieves comparable performance to the more common adjoint-based technique of topology optimization. While topology optimization enforces its fabrication constraints by subsequently applying filters over multiple optimization phases, MBSO allows for constraint functions to be defined explicitly from the parameterization and applied throughout a single-phase optimization. As a result, the shape-optimized structure converges much faster, using less than half as many sets of simulations as a topology-optimized structure (200 vs 471) — the computational requirements of both types of adjoint optimization are primarily limited by the number of simulations required. The shape-optimized structure also required an order of magnitude fewer parameters than the topology-optimized structure (1280 vs 40,401); although this did not strongly impact the computational performance in this case, it could result in significantly faster gradient calculations for larger-scale metasurfaces. However, while topology optimization requires very little knowledge or assumptions about what a final device might look like, MBSO does require an extra degree of judiciousness in choosing certain reasonable fixed parameters (in this case pillar pitch and height) as discussed in Section III.2.
@@
As an adjoint-based optimization strategy, MBSO’s scalability to larger geometrical areas is primarily limited by the computational requirements in memory and the time to perform the full-wave simulations at each iteration. Several techniques may be used to increase the viability of large-scale optimizations. Computational advancements such as GPU acceleration can enable larger-scale full-wave simulations [50]. Past authors have performed shape optimizations on structures an order of magnitude larger than the present work by applying symmetrical boundary conditions and using rectangular elements aligned to the FDTD mesh [34] or by using Maxwell solvers based on boundary integral methods [51, 52] — both strategies are particularly well-suited to shape optimizations. Other works have avoided full-wave simulations by using local periodic approximations [53] and coupled mode theory [54] in limited regimes or deep learning methods to predict forward and inverse problems [55, 56, 57]. The MBSO parameterization method could easily be adapted into any of these alternative optimization strategies.
@@
The MBSO methodology is readily applicable to the inverse design a variety of structures, especially when there is a need for specific shapes or distinct elements. Broadly, it can be used in lieu of topology optimization for any inverse design problem. In particular, MBSO is naturally suited for the inverse design of photonic structures such as photonic crystals, large-area metasurfaces, and diffraction gratings that are typically composed of arrays of simple shapes. Additionally, MBSO could be useful for designing reconfigurable metasurfaces — which have potential applications including imaging, beam steering for LiDAR, and displays —  such that the design space is limited to a fixed number of meta-atoms that can be independently addressed by some tuning mechanism [58]. For certain materials, it may be challenging to fabricate arbitrary shapes due to facet-selective etching or growth. In these instances, MBSO could be easily adapted to keep any edges aligned to particular crystal facets in order to optimize structures exactly how they would be fabricated, potentially enabling the inverse design of photonic structures onto new material platforms.
@@ Acknowledgements.
The authors thank Brian Edwards for fruitful conversations and Mathieu Ouellet for assistance with the three-dimensional image in Fig. 1(a). A.R.K. acknowledges support from the Fontaine Society and National Science Foundation Graduate Research Fellowship Program. A.R.K. and L.C.B. acknowledge support from the National Science Foundation under awards ECCS-1842655 and ECCS-2129183. N.E. acknowledges partial support from the Air Force Office of Scientific Research Multidisciplinary University Research Initiative under grant FA9550-21-1-0312.